OpenStudy (anonymous):

Calculate the inverse function and the domain of the functions: f(x)= 3 +ln(e^x- 2) So far this is my working. I'm not sure of what to do with the -2 y=3+ln(e^x-2) y-3=ln(e^x-2)

OpenStudy (anonymous):

so far this is my working f(x)= 3 +ln(e^x- 2) y=3+ln(e^x-2) y-3=ln(e^x-2) e^(y-3)=e^x-2 2+e^(y-3)=e^x 2+y-3=ln(e^x ) y-1=x y=x+1

OpenStudy (anonymous):

which grade are you?

OpenStudy (anonymous):

let h(x) = ln (e^x -2) g(x) = x +3 so that, you have f(x) = g(h(x) --> \(f^{-1}(x) = (g(h(x))^{-1} = h^{-1}(g^{-1}(x))\) Now, find out \(g^{-1}(x) = x-3\) replace into the expression above, you have \(f^{-1}(x) = h^{-1}(x-3)\) finding \(h^{-1}(x-3)\)

OpenStudy (anonymous):

\(h(x) = ln (e^x +2)\\h(x-3) = ln (e^{x-3}+2)\) inverse of this guy is the required function.

OpenStudy (anonymous):

Apply "switch the role" rule, you have \(e^{x-3} = ln (h^{-1}(x-3) +2)\) e both sides.... can you step up?? This problem is not a piece of cake, that's why I asked "what grade are you?". Hopefully, you are a senior in university to get what I mean.

OpenStudy (anonymous):

I'm first year university. This problem has me going. Thanks for the help :)